OpenStudy (anonymous):

people I do not get how to do this problem. x^2y^2+x^3 y. I am factoring the biggest polynomials.

OpenStudy (harsimran_hs4):

x^2y(y +x) is the required factorization (clear?) this is self explanatory though

OpenStudy (anonymous):

I do not get it my problem is x^2y^2+x^3y

OpenStudy (harsimran_hs4):

this is equivalent to xxyy + xyyy = xyy(x+y) = xy^2(x+y)

OpenStudy (anonymous):

would the answer be x(xyy+xxy) just kinda guessing.

OpenStudy (harsimran_hs4):

in this you can still take out xy common because both of them contain xy

OpenStudy (harsimran_hs4):

got it or still struck up?